Today I realized that the SQZ PLLs where no more locked after this intervention. I found the locking point of MAIN anc CC PLLs, for SC I have the impression that the laser switched off again (to be checked)
MAIN PLL: was moved from 25773 to 23800
CC PLL: was moved from 25820 to 23825
I managed to lock the CC PLL for a 10 mins and then it never locked again (to be checked with a spectrum analyzer). Also the main PLL is a way more instable than usual. It continue to unlock. I have the impression that we are working in multimode regime but it does not seems from SHG scan. The power of TEM00 mode is decreased in the new config (fg 1)
Considering the calbration of the MAIN PLL Peltier, we know that with 65536 bit we can do 16.666V this means that too coover the 1.4K done with PSL we should pass from 25800 to 31400. Around that point the main PLL status pass trough an high state but I don't manage to lock the PLL. Check with a spectrum analyzer.
The check of the squeezer should be possible by putting SQZ MAIN in DOWN then restore the old setpoint of both PLL in the GUI, switch off the MAIN PLL Fast loop and lock by hand the CC PLL.
But this configuration prevents all the injection in the interferometer.
To solve the situation it is worth a check with a spectrum analyzer in DET Eroom and after that understand if we can work on this point, otherwise the PSL temperature should be moved in case of one of the SQZ laser is in a multimode configuration